OpenStudy (anonymous):

2y-4x=8 (-2,8) decide wheather the given ordered pair is a solution of the eqation

OpenStudy (id21):

Is (-2,8) a solution? If x=-2 and y=8 Just put it in equation: \[2 \times 8-4\times(-2)=8?\] Is it true? If it is true, then (-2,8) is solution of this equation. 16+8=8??? It's not!!!! :D so this point (-2,8) is not a solution of this equation.